DIRC Operations Weekly Report
Period: Wednesday 2007/4/17 - Wednesday 2007/4/24
Oncall: Alejandro Perez
Fast Monitoring: Alejandro Perez, Bradley Wogsland
PEP peak lumi: ~1.0 x 1034
(previous week: ~1.0 x 1034)
BaBar integrated lumi: 2.13 fb-1
(previous week: 3.28 fb-1)
Run 6 PEP delivery: 2.19 fb-1
(previous week: 3.43 fb-1)
Run 6 total delivery: ~30.61 fb-1
Run 6 total BaBar recorded: 29.38 fb-1
including 0 fb-1 off-peak data
PEP activities this week:
4/18: in the morning 2 hours MD for beam characterization.
4/19: 12 hours ROD.
4/20 - 4/23: frequent MD studies at low luminosity.
4/24: Vacuum activity in HER. Suspect that VacSealing made on 4/9 (last 3 days ROD) near Q4 is not holding.
- Brace installed in the support tube, this reduced beam vertical oscillation by factor of 10.
- Z dumper installed between support tube and EMC.
- Accelerometers istalled on support tube and near Q4.
- 12 Decker skew quads istalled.
BaBar activities/issues this week:
4/19: 12 hours ROD.
- 4/19: 12 hours ROD,
- Ray put some stickers in the DIRC chiller saying the voltage of power supply (safety reasons).
- Nicolas put an alarm in the ALH about the power DIRC chiller glitches, with instructions to the pilot to tell the DQM to go check the DIRC chiller status.
- Chris O'Grady swapped a ROM with a repaired one in order to test it.
- Got paged at 4am due to SynchRequest error from the ROM Chris O'Grady swapped. The partition was rebooted and solved the problem.
- 4/20: In the morning we had optical fiber disconnection from the same ROM. So Chris replaced for the old one.
- 4/23: Got paged twice in the afternoon because of long periodes of high scaler rates. Told the pilot to be persistent in complaining MCC to do something about it.
logbook for details
- Replace two DFBs, one in sector 0 and one in sector 1 with a group of 8 innefficient or
flaky channels (require bridge).
- Replace new DFB in sector 8 with 18 (TDC chip 0) often dead channels and figure
out what is killing DFBs in that slot.
- Found a mini X-tree in sector 7 HV group 55 than only shows up during taking cosmics.
Jose, Alejandro and Bradley rotating backup & on-call commissioning